# Green Operations of SWIPT Networks: The Role of End-User Devices

### v1

@@ VI-B Linear Energy Harvesting
Fig. 2 suggests also that our modeling approach is very accurate across very diverse system configurations. In particular, very good accuracy is achieved even for low values of user density, for which the dense IoT regime assumption of Theorem 2 does not hold.
@@
These results suggest that, in sharp contrast to RANs delivering only connectivity, as the density of IoT users increases, their contribution to the service capacity of the whole SWIPT network (in the form of power delivered to other IoT users) increases too. The effect is visible in all configurations when comparing the setting in which the energy harvested originates only from BS transmissions (the ”no UPC” curves in Fig. 2), with those in which users can harvest energy also from transmissions of other users. In all configurations, we find that as user density increases, its effects manifest as a decrease in the power consumed by the network. Such a decrease continues until a minimum is reached, after which the power consumed increases again with increasing user density. Such behavior is quite surprising, but it is the result of the interplay between, on one side, the detrimental effect of rising interference on communications (which decreases the efficiency with which network resources are used), and the beneficial effect of the increase in user transmissions on the amount of power harvested by IoT users. Indeed, when increasing user density from very low values, these two contrasting effects cause the overall power consumed by the network to decrease, before increasing again due to the effects of high levels of interference on user-perceived QoS for data. The impact of user contribution to network service capacity is less marked when the target minimum service capacity is low (e.g., in the $1$ mW plots), due to the low impact of power delivery on the overall consumed power of the network, and it increases for higher target minimum harvested power. Such impact is limited also in the SPS mode, due to the lower efficiency of passive EH in that configuration.
@@
Figure 4: Increase in consumed power with respect to linear EH, as a function of user density, for HLP $1$ mW target minimum harvested power, and different IoT UE receiver configurations.
These features of optimal power consumption suggest that (possibly dynamic) configuration tuning holds the potential to play a key role in minimizing the energy footprint of a SWIPT network. Indeed, these features suggest that by taking advantage of user-provided service capacity, it is possible to make the network increasingly more energy efficient when the demand (of both energy and communication) it must satisfy grows. This is suggested also by Fig. 4, which shows indeed that for all configurations and target minimum harvested power, the network energy efficiency increases when user density increases. Here too, when the user-provided service capacity dominates the overall energy footprint of the network, the decrease of per-user energy consumed is faster, while it tends to slow down for those user densities for which the effects of interference dominate. Quite surprisingly, for the same target minimum harvested power, when user-provided service capacity dominates, the power per user consumed by the network does not depend on the amount of load proportionality of the BS energy model (red and blue curves are very close). It is only when interference dominates that the higher BS densities required to satisfy the target QoS for information transfer bring an increase in the impact of the fixed energy costs of BSs on the overall energy footprint of the network.
@@
Another noteworthy feature of a SWIPT network, emerging from Fig. 2 is that for each configuration of EH receiver architecture and target minimum harvested power, there is a value of user density below which no feasible solution of Problem 1 exists. This is due to the lower bound to the mean number of users per base stations (equal to $5$ in our evaluations). Indeed, as with increasing distance from the serving base station the power received by users decays much faster than channel capacity, guaranteeing a target QoS for harvested power at low user densities requires ever increasing BS densities and thus very low mean number of users per BS, up to the point of economic unfeasibility. Note however that this is relative to a scenario in which users are distributed uniformly at random in space. In realistic settings, this implies that deployments of small amounts of EH IoT devices are of course feasible, but they should be, at least locally, dense enough to be able to deliver power with the target QoS while satisfying the system constraints.
@@
(a) Time Switching
(b) Dynamic Power Splitting
(c) Static Power Splitting
(a) Time Switching
(b) Dynamic Power Splitting
(c) Static Power Splitting
@@
In Fig. 7, 7, and 7 we plotted the optimal values of BS density, of transmit power, and of time (respectively, power) split ratio, for the three EH receiver configurations. As Fig. 7 shows, in the TS and DPS cases the evolution of the optimal BS density as a function of user density is characterized by three regimes. At first, for very low user densities, the optimal BS density increases with user density as expected. For higher user densities, the optimal BS density reaches a plateau in the case without user-supplied service capacity, and it decreases in the case in which the full potential of passive EH is available. This suggests (quite surprisingly!) patterns of BS sleep modes which, in contrast to those for non-SWIPT RANs, turn off BSs when demand increases. Finally, for even larger user densities, in the interference-dominated regime, the optimal BS density increases again. These three regimes are not present however for all values of target minimum harvested power, due to the minimum user-per-BS limit adopted. These three regimes are not present in the SPS case, in which the lower efficiency of passive EH brings optimal BS densities which increase monotonically with user density.
From Fig. 7 we see that the optimal transmit power varies with user density in a specular manner to optimal BS density, decreasing when it increases, and vice versa. That is, the energy optimal strategies emerging from our GA algorithm tend to compensate for the thinning (resp. densification) of BS density with cell zooming (resp. shrinking), via transmit power tuning.
@@
VI-C Impact of harvesting nonlinearities
A crucial aspect of the energy harvesting process is the presence of nonlinearities in the relation between the output power of the energy harvester circuit at the IoT devices and its input power. Indeed, such aspect may affect heavily the amount of harvested power as a function of the network settings, and thus the energy optimal configuration of the whole SWIPT network. To investigate its impact on the energy optimal strategies derived by our GA, in another set of experiments we have considered the energy harvesting curve given by Equation 1. Following indications from the literature [30], we have set the maximum output power $h_{max}$ to $10$ mW, $\chi$ to $274$, and the sensitivity threshold $h_{s}$ to $0.064$ mW. Figure 9 shows the nonlinear EH curve for the chosen parameters and for two values of $\iota$, which tunes the maximum efficiency of the nonlinear harvesting process. Its values have been chosen not to have a maximum efficiency larger than one, and to approximate the linear harvesting curve which we have considered in the previous experiments. In the figure indeed, we have also plotted the linear harvesting functions whose slope approximates well that of the two sigmoids. As the plot shows, both sigmoids are characterized by a range of input power within which they are well approximated by a linear function (and which we denote henceforth as a ”quasilinear regime”). Note that, with the given settings, choosing $\iota=0.9$ (respectively, $1.30$) produces a sigmoid which in the quasilinear regime approximates well a linear model with slope $0.9$ (respectively, $0.77$).
@@ VI-C Impact of harvesting nonlinearities
Figure 9 illustrates the increase in power consumed at the optimum induced by the nonlinear EH function versus that of its linear approximation, for the given values of $\iota$, and for time splitting EH mode. As the plot suggests, nonlinearity brings to an increase of the consumed power for target minimum harvested power values which are much lower than $h_{max}$, and for low values of user density. Indeed, for a target harvested power of $1$ mW the sigmoid is in a regime in which the conversion efficiency is $20\%$ to $30\%$ less than its maximum value, pushing the system to compensate with a more conservative transmit power allocation, thus bringing to a higher overall consumed power. This is more evident for lower user densities, as in those regimes the lower mean amount of users per base station brings (by the law of small numbers) to larger variance and thus to greatly amplify the effect of the $20\%$ to $30\%$ decrease in conversion efficiency on those users whose harvested power falls on the tail of the distribution.

### v3

@@ VI-C Linear Energy Harvesting
Fig. 3 suggests also that our modeling approach is extremely accurate across very diverse system configurations. Indeed, the markers reporting simulation results overlap the curves obtained with our analytical model. In particular, very good accuracy is achieved even for low values of user density, for which the dense IoT regime assumption of Theorem 2 does not hold.
@@
These results suggest that, in sharp contrast to RANs delivering only connectivity, as the density of IoT users increases, their contribution to the service capacity of the whole SWIPT network (in the form of power delivered to other IoT users) increases too. The effect is visible in all configurations when comparing the setting in which the energy harvested originates only from BS transmissions (the ”no UPC” curves in Fig. 3), with those in which users can harvest energy also from transmissions of other users. In all configurations, we find that as user density increases, its effects manifest as a decrease in the power consumed by the network. Such a decrease continues until a minimum is reached, after which the power consumed increases again with increasing user density. Such behavior is quite surprising, but is the result of the interplay between, on the one side, the detrimental effect of rising interference on communications (which decreases the efficiency with which network resources are used), and on the other, the beneficial effect of the increase in user transmissions on the amount of power harvested by IoT users. When increasing user density from very low values, these two contrasting effects cause the overall power consumed by the network to decrease, before increasing again due to the effects of high levels of interference on user-perceived QoS for data. The impact of user contribution to network service capacity is less marked when the target minimum service capacity is low (e.g., in the $1$ mW plots), due to the low impact of power delivery on the overall consumed power of the network, and it increases for higher target minimum harvested power. Such impact is limited also in the SPS mode, due to the lower efficiency of passive EH in that configuration.
@@
Figure 5: Increase in the power per $km^{2}$ consumed by the network with respect to linear EH, as a function of user density, for HLP $1$ mW target minimum harvested power, and for different IoT UE receiver configurations.
These features of optimal power consumption suggest that (possibly dynamic) configuration tuning holds the potential to play a key role in minimizing the energy footprint of a SWIPT network. Indeed, these features suggest that by taking advantage of user-provided service capacity, it is possible to make the network increasingly more energy efficient when the demand (of both energy and communication) it must satisfy grows. This is suggested also by Fig. 5, which shows that for all configurations and target minimum harvested power, the network energy efficiency increases when user density increases. Here too, when the user-provided service capacity dominates the overall energy footprint of the network, the decrease of per-user energy consumed is faster, while it tends to slow down for those user densities for which the effects of interference dominate. Quite surprisingly, for the same target minimum harvested power, when user-provided service capacity dominates, the power per user consumed by the network does not depend on the amount of load proportionality of the BS energy model (red and blue curves are very close). It is only when interference dominates that the higher BS densities required to satisfy the target QoS for information transfer bring an increase in the impact of the fixed energy costs of BSs on the overall energy footprint of the network.
@@
Another noteworthy feature of a SWIPT network emerging from Fig. 3 is that for each configuration of EH receiver architecture and target minimum harvested power, there is a value of user density below which no feasible solution of Problem 1 exists. This is due to the lower bound to the mean number of users per base stations (equal to $5$ in our evaluations). Indeed, as with increasing distance from the serving base station the power received by users decays much faster than channel capacity, guaranteeing a target QoS for harvested power at low user densities requires ever increasing BS densities and thus very low mean number of users per BS, up to the point of economic unfeasibility. Note however that this is relative to a scenario in which users are distributed uniformly at random in space. In realistic settings, this implies that deployments of small amounts of EH IoT devices are of course feasible, but they should be, at least locally, dense enough to be able to deliver power with the target QoS while satisfying the system constraints.
(a) Time Switching
(b) Dynamic Power Splitting
(c) Static Power Splitting
@@
(a) Time Switching
(b) Dynamic Power Splitting
(c) Static Power Splitting
@@
In Fig. 8, 8, and 8 we plot the optimal values of BS density, of transmit power, and of time split ratio (respectively, power split ratio), for the three EH receiver configurations. As Fig. 8 shows, in the TS and DPS cases the evolution of the optimal BS density as a function of user density is characterized by three regimes. First, for very low user densities, the optimal BS density increases with user density, as expected. For higher user densities, the optimal BS density reaches a plateau in the case without user-supplied service capacity, and it decreases in the case in which the full potential of passive EH is available. This feature suggests new and quite surprising patterns of BS sleep modes which, in contrast to those for non-SWIPT RANs, turn off BSs when demand increases. Finally, for even larger user densities, in the interference-dominated regime, the optimal BS density increases again. These three regimes are not present however for all values of target minimum harvested power, due to the minimum user-per-BS limit adopted. These three regimes are not present in the SPS case, in which the lower efficiency of passive EH brings optimal BS densities which increase monotonically with user density.
From Fig. 8 we see that the optimal transmit power varies with user density in a specular manner to optimal BS density, decreasing when it increases, and vice versa. That is, the energy optimal strategies emerging from our GA algorithm tend to compensate for the thinning (resp. densification) of BS density with cell zooming (resp. shrinking), via transmit power tuning. This is likely due to the fact that, when IoT node density increases, the impact on the system of the inefficiency due to high distance between BS and EH IoT user is stronger, pushing the optimal solution towards larger BS densities.
@@
Finally, to characterize the effectiveness of our GA algorithm, we have compared the results derived from it with those achieved with a simple grid search heuristic. This latter is based on the discretization of the optimization parameters, and on an exhaustive search over these discrete values. In all tests, over all considered settings and for different values of the discretization step, the difference in results were only due to the chosen discretization step, and could be made arbitrarily small with a finer discretization step. This suggests that, for all the settings we have considered, our GA algorithm proved effective in avoiding the search process to get stuck into local minima, while being substantially more efficient.
VI-D Impact of harvesting nonlinearities
A crucial aspect of the energy harvesting process is the presence of nonlinearities in the relation between the output power of the energy harvester circuit at the IoT devices and its input power. Indeed, such aspect may heavily affect the amount of harvested power as a function of the network settings, and thus the energy-optimal configuration of the whole SWIPT network. To investigate its impact on the energy-optimal strategies derived by our GA, in another set of experiments we considered the energy harvesting curve given by Equation 1. Following indications from the literature [30], we have set the maximum output power $h_{max}$ to $10$ mW, $\chi$ to $274$, and the sensitivity threshold $h_{s}$ to $0.064$ mW. Figure 10 shows the nonlinear EH curve for the chosen parameters and for two values of $\iota$, which tunes the maximum efficiency of the nonlinear harvesting process. Its values have been chosen not to have a maximum efficiency larger than one, and to approximate the linear harvesting curve which we have considered in the previous experiments. In the figure, we have also plotted the linear harvesting functions whose slope approximates well that of the two sigmoids. As the plot shows, both sigmoids are characterized by a range of input power within which they are well approximated by a linear function (and which we denote henceforth as a ”quasilinear regime”). Note that, with the given settings, choosing $\iota=0.9$ (respectively, $1.30$) produces a sigmoid which in the quasilinear regime approximates well a linear model with slope $0.9$ (respectively, $0.77$).
@@ VI-D Impact of harvesting nonlinearities
Figure 10 illustrates the increase in power consumed at the optimum induced by the nonlinear EH function versus that of its linear approximation, for the given values of $\iota$, and for time splitting EH mode. As the plot suggests, nonlinearity brings to an increase of the consumed power for target minimum harvested power values which are much lower than $h_{max}$, and for low values of user density. For a target harvested power of $1$ mW the sigmoid is in a regime in which the conversion efficiency is $20\%$ to $30\%$ less than its maximum value, pushing the system to compensate with a more conservative transmit power allocation, thus bringing to a higher overall consumed power. This is more evident for lower user densities, as in those regimes the lower mean amount of users per base station brings (by the law of small numbers) to larger variance and thus to greatly amplify the effect of the $20\%$ to $30\%$ decrease in conversion efficiency on those users whose harvested power falls on the tail of the distribution.